Meeting Minutes

20 July 2004

In attendance: J. Nelson (notetaker), R. Hall, J. Rock, D. Rogind

 

Structure Summary Display

 

SLED FDBK through SCP slow feedback

†††† Thereís a bug in Janiceís code that she needs to fix.Comm:delta doesnítí go to zero everytime it should

 

paoutdata mat files compatible with matlab 7?

†††† Done deal.Yeah, Bob!

††††

MacNair module support of aes

(Janice didnít do this)Janice will ask Doug about signals into LOTS of channels.

Letís get a real acquire button first Ė that triggers all the MacNairs.Put the button on a tr06 module panel

(still true, not needed till September) Ron will work with Bob to get the gain in pk-pk volts archived as a float.

 

2-pack PLC coming in August-ish.

Janice reports: no money till October 1st.

Specification avalanche has occurred.See:

http://www.slac.stanford.edu/~jnelson/jln_todo_xls.htm 2 pack section

Estimated installation of mod + klys in August, RF in early September

 

NLCTA/8-pack summary display

Janice fixed this. Stn2 gradients get stuck when the power goes to zero.

Save/restore of trip:ring Ö letís convert the sequence to waveform (JN/RC)

 

AFG driver code is done

Debbie and Janice are gonna talk after to discuss whatís next

 

Matlab/epics 8-pack SLED feedback + A/P from I/Q calc Ö

Done.Epicsification is yet to be done.

(still true) Janice will work with Ron to find common analysis

 

HamHands for NLCTA

Janice sent out a spec for 6 SISes in TR00, 3 for each of stations 1 & 2.Need acquires for these two triplets plus one for MacNair modules.Once acquires are in, will want relevant processing loops to hit acquire buttons.

(Janice) Reset trips, elog signals/status, more SISes

Janice thinks this is desired within the next few months.

 

(done) Vacuum- Almost ready HamHandsNlcta watches for vacuum trips of Stns 1 or 2, makes an elog entry, and resets the station!!!

 

CMLOG clogging

Judy gonna bug James to think about it.TR00 is the only one having these problems now and it seems to be independent of reboots.

 

TR08 CPU maxing out

(still true) Ron suspects low priority callback routine.Proc Loop is whatís using this.We could try getting rid of some of the Pave/min/max calcs for uncritical powers.

 

SCP button to call pre-configed archive browser

This is done.

We will have this capability from the Unix Foyer when it is released soon.

 

Auto-elogging of archived data

(This was horribly complicated.)

 

Reboot IOCs with new directory structure

(still true)TR00/01 coming soon.

 

Realm split news

Judy reported that realm split may happen in the next couple weeks.For us, this means we get our own cmlog (without pep2 epics messages), cw+alh move from gtw02 to gtw04.